OpenStudy (anonymous):

Julie is a math tutor in the academic center. She charges $210 for the first 7 hours of tutoring and $20 for each additional hour of tutoring. If she earned $410 tutoring 1 student, how many total hours did she spent tutoring the student? Which of these equations represents this problem

OpenStudy (lgbasallote):

do you have any thoughts?

OpenStudy (anonymous):

7=410-210/20

OpenStudy (lgbasallote):

hmm not exactly

OpenStudy (lgbasallote):

if you let x be the additional hours..then 20x will be the cost of the extra hours right?

OpenStudy (anonymous):

A. h = 7 × 410 + 210 20 B. h = 7 − 410 + 210 20 C. h = 7 − 410 − 210 20 D. h = 7 + 410 − 210 20

OpenStudy (lgbasallote):

what are those 20?

OpenStudy (anonymous):

yea divide 20

OpenStudy (lgbasallote):

i cannot imagine what your options look like so just follow what im saying...do you agree with what i said earlier on?

OpenStudy (lgbasallote):

20x is the cost of extra hours?